OpenStudy (anonymous):

Can someone break down the 5 rules of logarithms for me?

OpenStudy (anonymous):

\[\log{a}+\log b=\log(ab)\] \[\log a- logb=\log(a/b)\] \[m \log a=\log a^m\] \[\frac{\log a}{\log b}=\log_{b}a\] \[x^{\log_xa}\]

OpenStudy (perl):

you need a base c for your change of basis formula log_b a = log_c a / log_c b

OpenStudy (lgbasallote):

^actually...NotSObright is right...:)

OpenStudy (perl):

i didnt say its wrong, but it can be another base, than the default base

OpenStudy (anonymous):

i only remember power/ quotient/ and power rules was there more to it?

OpenStudy (anonymous):

product

OpenStudy (perl):

see above

OpenStudy (anonymous):

In all the above formulas is for all bases ob default valid until you use that same base

OpenStudy (perl):

all bases valid , yes

OpenStudy (anonymous):

so rules only apply or can be different under certain circumstances?

OpenStudy (lgbasallote):

depends on the situation

OpenStudy (anonymous):

You learn through practice

OpenStudy (perl):

ok here are some examples log 2 + log 3 = log (2*3). ln (2/3) = ln 2 - ln 3

OpenStudy (perl):

log 2 + ln 3 does not combine, because they are different bases

OpenStudy (perl):

log 2^3 = 3* log 2 , see if you can find which rule I used

OpenStudy (perl):

so log 8 = log (2^3) = 3 log 2 so log 8 = 3 log 2

OpenStudy (anonymous):

ok it cleared it up a bit ty m8

OpenStudy (anonymous):

rule of power right?

Directrix (directrix):

Summary of log laws and demo problems at the link ---> http://www.themathpage.com/aprecalc/logarithms.htm#laws